Solve the given equation by Gauss elimination method
x₁ + 2x2-3x3 = -1
-3x1 + x2-9x3 = -7
5x+3x2 - 4x3 = 2

Answer by Alan3354(69443)

Solve the given equation by Gauss elimination method
---
Not "an equation" it's a system of equations.
---
Using x, y & z would be better, IMO.
=====================================
x₁ + 2x2-3x3 = -1
-3x1 + x2-9x3 = -7
5x+3x2 - 4x3 = 2